A packet of potatoes costs R8 more than a bag of tomatoes .A bag of onions costs R2 less than the bag of tomatoes .
If 3 packets of potatoes , 5 bags of onions and 4 bags of tomatoes cost R350.60, calculate the individual prices of these items .

Answer by richwmiller(17219)

p=t+8
o=t-2
3p+5o+4t=350.6
3(t+8)+5(t-2)+4t=350.6
3t+24+5t-10+4t=350.60
12t+14=350.60
12t=336.60
t=28.05
p=36.05
o=26.05
check
3(36.05)+5(26.05) +4(28.05)=350.6
ok